Half of Henry's age added to 1/3 od Daisy's age is 11 years.
Six years from now, the sum of their ages will be 40 years.
How old is each?
: Half of Henry's age added to 1/3 od Daisy's age is 11 years.
Six years from now, the sum of their ages will be 40 years.
How old is each?

Answer by checkley77(3423)

.5H+D/3=11
(H+6)+(D+6)=40
H+D+12=40
H+D=40-12
H+D=28
H=28-D NOW REPLACE H BY (28-D) IN THE FIRST EQUATION
.5(28-D)+D/3=11
14-.5D+D/3=11
-.5D+D/3=11-14
(3*-.5D+D)/3=-3
(-1.5D+D)/3=-3
-.5D=-9
D=-9/-.5
D=18 ANSWER FOR DAISY'S AGE NOW.
H+18+12=40
H+30=40
H=40-30
H=10 AGE OF HENRY NOW.
PROOF:
.5*10+18/3=11
5+6=11
11=11